Cayley theorem for monoidoids Gheorghe Ivan, Timişoara, Romania

Keywords: monoidoid, monoid bundle, mappings monoidoid, quasipermutation groupoid AMS Subject classification: 20L05, 20L99 Abstract: In this paper we define the monoidoids and give a brief summary of their most important properties. The main result is the Cayley’s theorem for monoidoids.